# WISE: A World Knowledge-Informed Semantic Evaluation for Text-to-Image Generation

### v2

@@ Abstract
Text-to-Image (T2I) models are capable of generating high-quality artistic creations and visual content. However, existing research and evaluation standards predominantly focus on image realism and shallow text-image alignment, lacking a comprehensive assessment of complex semantic understanding and world knowledge integration in text-to-image generation. To address this challenge, we propose WISE, the first benchmark specifically designed for World Knowledge-Informed Semantic Evaluation. WISE moves beyond simple word-pixel mapping by challenging models with 1000 meticulously crafted prompts across 25 subdomains in cultural common sense, spatio-temporal reasoning, and natural science. To overcome the limitations of traditional CLIP metric, we introduce WiScore, a novel quantitative metric for assessing knowledge-image alignment. Through comprehensive testing of 22 models (10 dedicated T2I models and 12 unified multimodal models) using 1,000 structured prompts spanning 25 subdomains, our findings reveal significant limitations in their ability to effectively integrate and apply world knowledge during image generation, highlighting critical pathways for enhancing knowledge incorporation and application in next-generation T2I models. Code and data are available at PKU-YuanGroup/WISE.
@@ 1 Introduction
Text-to-image (T2I) models [53] are able to generate high-quality images that visually match explicit text descriptions. However, these models often have difficulty in ensuring true factual accuracy, especially when dealing with prompts that require complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching). This shortcoming mainly stems from their limited ability to model the information behind world knowledge [51] - this information constitutes the massive facts and complex relationships required for real-world understanding. Although current unified multimodal models have begun to use LLMs’ powerful text modeling and implicit information extraction capabilities to break through this bottleneck through specific representations combined with generative model decoders [28; 41], existing evaluation benchmarks still have obvious limitations. They only focus on the surface alignment of images and texts, and fail to evaluate the core capabilities of T2I models from the perspective of implicit understanding and intrinsic knowledge matching, which seriously hinders the development of intelligent T2I systems.
@@
Figure 1: Comparison of previous straightforward benchmarks and our proposed WISE. (a) Previous benchmarks typically use simple prompts, such as "A photo of two bananas" in GenEval [9], which only require shallow text-image alignment. (b) WISE, in contrast, uses prompts that demand world knowledge and reasoning, such as "Einstein’s favorite musical instrument," to evaluate a model’s ability to generate images based on deeper understanding.
@@
Specifically, most T2I benchmarks suffer from a lack of semantic complexity. As shown in Figure 1, they use overly straightforward and simple prompts, failing to effectively challenge models’ ability to understand and generate images based on the model’s world knowledge. Furthermore, the most commonly used metric FID [12], primarily focuses on the realism of generated images. Although some benchmarks [11; 46; 50] utilize models like CLIP [31] to assess image-text semantic consistency scoring. However, CLIP’s limitations [52] in capturing fine-grained semantic information and handling complex reasoning hinder a comprehensive assessment of models’ performance in processing intricate semantic information. Consequently, existing evaluations fail to fully reveal the potential of models in real-world scenarios, particularly in tasks requiring world knowledge. For instance, when generating an image depicting a "tadpole that has undergone metamorphosis," a model needs not only to comprehend the textual description ("tadpole", "metamorphosis") but also to invoke its internal world knowledge. This includes understanding amphibian development, the specific morphological changes involved (e.g., the growth of legs, the loss of the tail, the development of lungs), and the biological processes driving this transformation.
@@
We broadened the scope of our evaluation beyond traditional dedicated T2I models. We employed our novel benchmark, WISE, to evaluate a total of 22 T2I models, encompassing both 10 dedicated T2I models and 12 unified multimodal models. However, experiment results demonstrate significant deficiencies in complex semantic understanding and world knowledge integration across existing T2I models. Even for unified multimodal models, their strong understanding capabilities do not fully translate into advantages in image generation, as revealed by our WISE evaluation. Despite their theoretical advantages, unified models generally underperform compared to dedicated T2I models in leveraging world knowledge for image generation. This indicates that current approaches to integrating LLMs within unified multimodal models may not yet fully unlock their potential for image generation that effectively integrates and applies world knowledge. Our main contributions are as follows:
@@
  • Our experiment results firstly demonstrate that existing T2I models exhibit significant limitations in integrating world knowledge, while the unified model underperforms even compared to dedicated T2I models.
@@ 2.1 Dedicated T2I Models
Dedicated T2I models represent the mainstream approach in the T2I field and have achieved remarkable progress in recent research. Currently, these models primarily fall into two categories: autoregressive models and diffusion models. Autoregressive [3; 6; 10; 40; 37; 29] models treat image generation as a sequence generation problem, similar to text generation. However, due to their computational cost and limitations in image quality, diffusion models have become the dominant paradigm. Diffusion [13; 32; 18; 34; 23] models iteratively add noise to an image and then progressively denoise it, often using pre-trained text encoders (e.g., CLIP [31]) to transform text prompts into embeddings that guide the denoising process. Key advancements include GLIDE [27], which pioneered diffusion models for T2I; Latent Diffusion Models (LDMs) [33], which improve quality and efficiency by operating in latent space; and Stable Diffusion series [30; 5], a landmark achievement built on LDMs.
@@ 2.2 Unified Multimodal Models
Unified multimodal models aim to construct general-purpose models capable of processing both textual and visual inputs, and performing cross-modal generation and understanding. These models [39; 16; 8; 43; 44; 25; 4; 48; 17; 45; 21; 38; 41; 35] are typically built upon powerful large language models (LLMs) [54] and extend next-token prediction [2] to image generation: the LLM generates visual tokens, and a VQ-VAE [42] or Diffusion model serves as a detokenizer. Moreover, Transfusion [55] and Show-O [49] demonstrate that bidirectional image diffusion can be combined with autoregressive text prediction within the same framework. D-DiT [22] achieves both Text-to-Image (T2I) and Image-to-Text (I2T) tasks using an end-to-end diffusion model. A crucial question concerning unified multimodal models is whether their understanding and generation capabilities can mutually enhance each other. Some studies [41; 45] have provided evidence supporting this phenomenon. However, in contrast to the rich and comprehensive benchmarks for multimodal understanding, T2I benchmarks are often relatively simple, lacking an in-depth examination of complex semantic understanding and world knowledge reasoning, making it difficult to fully prove the phenomenon.
@@ 2.3 Text to Image Evaluation
Despite the Fréchet Inception Distance (FID) [12] being one of the most widely adopted metrics for evaluating the quality of generated images, it falls short in assessing text-image consistency, thus failing to comprehensively measure the capabilities of text-to-image models. To address this deficiency, researchers have introduced a series of more sophisticated and challenging benchmarks [36; 24; 47] and evaluation metrics [11; 19; 24]. For instance, DPG-Bench [14] focuses on evaluating models’ ability in dense prompt following. T2I-CompBench [15] provides a benchmark suite for evaluating compositional generation, where prompts typically combine multiple distinct attributes. Furthermore, GenEval [9] serves as an object-centric evaluation framework specifically designed to assess compositional attributes of images, such as object co-occurrence, position, number, and color. However, the prompts used in these benchmarks are mostly straightforward, primarily examining models’ ability to follow compositional instructions for generation. Recently, some studies have shifted focus towards evaluating the application of specific types of knowledge in T2I models, such as physical reasoning in PhyBench [26] and broad common-sense knowledge in Commonsense-T2I [7]. Nevertheless, these emerging benchmarks are still very limited in their research scope and the quantity of evaluations.
@@ 3.2 Discovering the Deep Visual Language Alignment Bottlenecks of Traditional Evaluation
Existing metrics such as CLIP-Score [11] and VQA-Score [19] offer partial solutions for evaluating text-to-image alignment, but both fall short when faced with complex, knowledge-intensive prompts. CLIP-Score, constrained by CLIP’s bag-of-words limitation, struggles with compositional semantics and relational reasoning. VQA-Score improves on this by leveraging likelihood estimation via question answering, but still lacks sensitivity to implicit visual understanding and world knowledge grounding.To address these limitations, we propose a multi-faceted evaluation protocol to rigorously assess the quality of generated images, focusing on four key aspects: Consistency, Realism, Aesthetic Quality, and a composite metric, WiScore, shown in Figure 2.
Figure 4: Correlation between different metrics and human assessment.
@@
$\begin{gathered}\text{WiScore}=\frac{\alpha_{1}\times\text{Consistency}+\alpha_{2}\times\text{Realism}+\alpha_{3}\times\text{Aesthetic Quality}}{2},\\ \text{subject to }\alpha_{1}+\alpha_{2}+\alpha_{3}=1.\end{gathered}$ (1)
In this paper, we set $\alpha_{1}=0.7$, $\alpha_{2}=0.2$, and $\alpha_{3}=0.1$. This configuration not only prioritizes Consistency, reflecting the importance of accurately representing the prompt’s intended objects and their relationships, but also incorporates Realism and Aesthetic Quality to ensure overall image quality. Meanwhile, we observe that WiScore exhibits a strong positive correlation with human evaluations under certain hyperparameter settings, as detailed in Appendix B. A higher WiScore indicates superior performance in accurately depicting objects and concepts based on world knowledge. The detailed scoring criteria and complete results for each component score are provided in Appendix C. In the evaluation, we employ GPT-4o-2024-05-13, a powerful multimodal large language model, as the evaluator to assess the performance of T2I models. A carefully designed and rigorous scoring protocol is adopted to ensure consistency and reliability, with full implementation details provided in the Appendix D. As shown in Figure 4, among all automated metrics, WiScore exhibits the highest agreement with human judgments. The detailed comparison process can be found in the Appendix E.
@@ 4.1 Experiment Settings
We evaluate 22 T2I models, including 10 dedicated T2I models and 12 unified multimodal models. The dedicated T2I models are: stable-diffusion-v1-5 [33], stable-diffusion-2-1 [33], stable-diffusion-xl-base-0.9 [30], stable-diffusion-3-medium [5], stable-diffusion-3.5-medium [5], stable-diffusion-3.5-large [5], playground-v2.5-1024px-aesthetic [20], PixArt-XL-2-1024-MS [1], FLUX.1-dev [18], and FLUX.1-schnell [18]. The unified multimodal models are: GPT4o, Metaquery [28], Janus-Pro-7B [4], Janus-Pro-1B [4], JanusFlow-1.3B [25], Janus-1.3B [44], show-o [49], show-o-512 [49], Orthus-7B-instruct [17], vila-u-7b-256 [48], Liquid [45] and Emu3 [43]. For image generation, we use the official default configurations of each model and fix the random seed to ensure reproducibility. All experiments are conducted on eight NVIDIA A800 GPUs. For short, we use SD to denote stable-diffusion, playground-v2.5 for playground-v2.5-1024px-aesthetic, and PixArt-Alpha for PixArt-XL-2-1024-MS.
@@ 4.2 Main Results
Model Cultural Time Space Biology Physics Chemistry Overall
Dedicated T2I
@@
Unified MLLM
GPT4o 0.81 0.71 0.89 0.83 0.79 0.74 0.80
@@
JanusFlow-1.3B 0.13 0.26 0.28 0.20 0.19 0.11 0.18
Janus-Pro-1B 0.20 0.28 0.45 0.24 0.32 0.16 0.26
Janus-Pro-7B 0.30 0.37 0.49 0.36 0.42 0.26 0.35
Orthus-7B-instruct 0.23 0.31 0.38 0.28 0.31 0.20 0.27
show-o 0.28 0.36 0.40 0.23 0.33 0.22 0.30
show-o-512 0.28 0.40 0.48 0.30 0.46 0.30 0.35
@@
Understanding-Generation Gap in T2I Tasks.Table 1 presents the WiScore of various models on the WISE benchmark, designed to evaluate the extent to which models leverage world knowledge in image generation. Overall, the results indicate that the GPT4o model significantly outperforms all others across all categories and the overall score. This suggests its superior capabilities in utilizing complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching) for image generation. The remaining open-source models demonstrate comparatively lower performance, with Unified MLLMs generally lagging behind Dedicated T2I models. This highlights that while Unified MLLMs are trained on large-scale image-text pairs and vast textual datasets, thereby acquiring substantial text and image understanding capabilities, effectively integrating this knowledge into the image generation process remains a significant challenge. Despite the overall underperformance of Unified MLLMs, the MetaQuery-XL model exhibits a prominent performance, achieving the best score among the open-source models, showcasing the potential of unified models to apply world knowledge for image generation.
Cross-Dimensional Performance Statical Analysis. This part systematically examines the performance disparities between the two model categories across various cognitive dimensions, employing a suite of statistical methods including paired-sample t-tests, Wilcoxon signed-rank tests, and effect size analysis. The paired t-test results indicate that T2I models demonstrate a significantly higher average score in the first three dimensions (Cultural/Time/Space) ($0.56\pm 0.08$) compared to the latter three (Biology/Physics/Chemistry) ($0.43\pm 0.07$), $t(9)=5.47$, $p<0.001$, with an effect size of Cohen’s $d=1.73$. This suggests a highly statistically and practically significant difference. To verify the robustness of these findings, a non-parametric Wilcoxon signed-rank test was conducted, yielding consistent conclusions ($W=55$, $p=0.002$). Similarly, the Unified MLLM model group also exhibited a characteristic pattern where the first three dimensions ($0.80\pm 0.19$) significantly outperformed the latter three ($0.79\pm 0.19$) ($t(10)=3.76$, $p=0.003$, $d=1.09$; Wilcoxon $W=66$, $p=0.005$). A Mann-Whitney U test revealed that although the magnitude of inter-dimensional differences was significantly larger for T2I models ($\Delta=0.13$) compared to Unified MLLMs ($\Delta=0.01$) ($U=32$, $p=0.057$), all models displayed an advantage in cultural tasks. This pattern of differences may stem from: 1) the Cultural/Time/Space dimensions relying more heavily on visual representation capabilities, a known strength of T2I models; and 2) scientific tasks requiring stronger cross-modal reasoning abilities, in which MLLM models (such as GPT4o) exhibit more balanced performance ($\Delta=0.01$). Notably, within the SD series, from v1.5 to 3.5-large, the dimensional difference decreased from 0.08 to 0.05; likewise, in the Janus series, from 1.3B to Pro-7B, the difference diminished from 0.04 to 0.03. This suggests that newer generation models are improving in their ability to balance performance across different dimensions.
@@
Figure 5: WiScore on rewritten prompts of different models. These prompts were simplified from the original WISE benchmark using GPT-4o (e.g., "The plant often gifted on Mother’s Day" to "Carnation"). Green ball indicates score increase after rewriting; red ball indicates score decrease.
We conducted an additional experiment to further highlight the limitations of current T2I models in processing world knowledge and explore possible solutions. We rewrote the prompts in the WISE benchmark using GPT-4o-2024-05-13, transforming them from complex, knowledge-requiring prompts to straightforward prompts. For example, the original prompt "Common plants for Mother’s Day" was rewritten to "Carnation". The detailed instructions for prompt rewriting are detailed in Appendix F. The same WiScore evaluation method was used for this experiment, and the results are presented in  Figure 5.
Surprisingly, nearly all models demonstrated significant performance improvements when evaluated with the rewritten prompts. The average WiScore increase across all Dedicated T2I models was +0.22, while Unified MLLMs saw an even larger average increase of +0.27. This suggests that a significant portion of the difficulty in the original WISE benchmark stemmed from the models’ inability to effectively parse and contextualize complex, knowledge-intensive prompts. By simplifying the prompts, we effectively removed this bottleneck, allowing the models to better showcase their generative capabilities. Notably, the models with the largest overall improvements tended to be those with lower initial scores on the original benchmark, indicating that prompt complexity had a disproportionately negative impact on their performance. For example, Janus-Pro-1B demonstrated an overall WiScore improvement of +0.34 after rewriting, highlighting the sensitivity of weak models to prompt complexity.
However, even with this simplification, the scores of both T2I and unified models, while improved, did not reach a level indicative of a complete and satisfactory generation of world knowledge across all categories. Furthermore, it highlights the core issue of lower-quality generation results in current unified models. Even MetaQuery-XL experienced a decrease in its score for the Physics category. This demonstrates that solely relying on enhancing the model’s ability to interpret world knowledge remains constrained by the bottleneck of generative capacity. Concurrently, the substantial performance gap underscores the need for future research to focus on improving model training methods and architectures, rather than solely relying on tricks aimed at enhancing the understanding level, such as prompt engineering.

### v3

# WISE: World Knowledge-Informed Semantic Evaluation for Text-to-Image Generation

@@ Abstract
Text-to-Image (T2I) models are capable of generating high-quality artistic creations and visual content. However, existing research and evaluation standards predominantly focus on image realism and shallow text-image alignment, lacking a comprehensive assessment of complex semantic understanding and world knowledge integration in text-to-image generation. To address this challenge, we propose WISE, the first benchmark specifically designed for World Knowledge-Informed Semantic Evaluation. WISE moves beyond simple word-pixel mapping by challenging models with 1000 meticulously crafted prompts across 25 subdomains in cultural common sense, spatio-temporal reasoning, and natural science. To overcome the limitations of traditional CLIP metric, we introduce WiScore, a novel quantitative metric for assessing knowledge-image alignment. Through comprehensive testing of 20 models (10 dedicated T2I models and 10 unified multimodal models) using 1,000 structured prompts spanning 25 subdomains, our findings reveal significant limitations in their ability to effectively integrate and apply world knowledge during image generation, highlighting critical pathways for enhancing knowledge incorporation and application in next-generation T2I models. Code and data are available at PKU-YuanGroup/WISE.
@@ 1 Introduction
Text-to-image (T2I) models [zhang2023text] are able to generate high-quality images that visually match explicit text descriptions. However, these models often have difficulty in ensuring true factual accuracy, especially when dealing with prompts that require complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching). This shortcoming mainly stems from their limited ability to model the information behind world knowledge [yu2023kola] - this information constitutes the massive facts and complex relationships required for real-world understanding. Although current unified multimodal models have begun to use LLMs’ powerful text modeling and implicit information extraction capabilities to break through this bottleneck through specific representations combined with generative model decoders [pan2025transfer, tong2024metamorph], existing evaluation benchmarks still have obvious limitations. They only focus on the surface alignment of images and texts, and fail to evaluate the core capabilities of T2I models from the perspective of implicit understanding and intrinsic knowledge matching, which seriously hinders the development of intelligent T2I systems.
@@
Figure 1: Comparison of previous straightforward benchmarks and our proposed WISE. (a) Previous benchmarks typically use simple prompts, such as "A photo of two bananas" in GenEval [ghosh2024geneval], which only require shallow text-image alignment. (b) WISE, in contrast, uses prompts that demand world knowledge and reasoning, such as "Einstein’s favorite musical instrument," to evaluate a model’s ability to generate images based on deeper understanding.
@@
Specifically, most T2I benchmarks suffer from a lack of semantic complexity. As shown in Figure 1, they use overly straightforward and simple prompts, failing to effectively challenge models’ ability to understand and generate images based on the model’s world knowledge. Furthermore, the most commonly used metric FID [heusel2017gans], primarily focuses on the realism of generated images. Although some benchmarks [hessel2021clipscore, wu2023human, xu2024imagereward] utilize models like CLIP [radford2021learning] to assess image-text semantic consistency scoring. However, CLIP’s limitations [yuksekgonul2022and] in capturing fine-grained semantic information and handling complex reasoning hinder a comprehensive assessment of models’ performance in processing intricate semantic information. Consequently, existing evaluations fail to fully reveal the potential of models in real-world scenarios, particularly in tasks requiring world knowledge. For instance, when generating an image depicting a "tadpole that has undergone metamorphosis," a model needs not only to comprehend the textual description ("tadpole", "metamorphosis") but also to invoke its internal world knowledge. This includes understanding amphibian development, the specific morphological changes involved (e.g., the growth of legs, the loss of the tail, the development of lungs), and the biological processes driving this transformation.
@@
We broadened the scope of our evaluation beyond traditional dedicated T2I models. We employed our novel benchmark, WISE, to evaluate a total of 20 T2I models, encompassing both 10 dedicated T2I models and 10 unified multimodal models. However, experiment results demonstrate significant deficiencies in complex semantic understanding and world knowledge integration across existing T2I models. Even for unified multimodal models, their strong understanding capabilities do not fully translate into advantages in image generation, as revealed by our WISE evaluation. This indicates that current approaches to integrating LLMs within unified multimodal models may not yet fully unlock their potential for image generation that effectively integrates and applies world knowledge. Our main contributions are as follows:
@@
  • Our experiment results highlight significant limitations in current T2I models’ ability to effectively integrate and apply world knowledge during image generation, revealing the potential directions for the fueture model optimization.
@@ 2.1 Dedicated T2I Models
Dedicated T2I models represent the mainstream approach in the T2I field and have achieved remarkable progress in recent research. Currently, these models primarily fall into two categories: autoregressive models and diffusion models. Autoregressive [chen2020generative, fan2024fluid, han2024infinity, tian2024visual, sun2024autoregressive] models treat image generation as a sequence generation problem, similar to text generation. However, due to their computational cost and limitations in image quality, diffusion models have become the dominant paradigm. Diffusion [ho2020denoising, ramesh2022hierarchical, flux2024, saharia2022photorealistic] models iteratively add noise to an image and then progressively denoise it, often using pre-trained text encoders (e.g., CLIP [radford2021learning]) to transform text prompts into embeddings that guide the denoising process. Key advancements include GLIDE [nichol2021glide], which pioneered diffusion models for T2I; Latent Diffusion Models (LDMs) [rombach2022high], which improve quality and efficiency by operating in latent space; and Stable Diffusion series [podell2023sdxl, esser2024scaling], a landmark achievement built on LDMs.
@@ 2.2 Unified Multimodal Models
Unified multimodal models aim to construct general-purpose models capable of processing both textual and visual inputs, and performing cross-modal generation and understanding. These models [team2024chameleon, jin2023unified, ge2024seed, wang2024emu3, wu2024janus, ma2024janusflow, chen2025janus, wu2024vila, kou2024orthus, wu2024liquid, li2024synergen, sun2024multimodal, tong2024metamorph, shi2024llamafusion, deng2025emerging, lin2025uniworld, li2025uniworld, li2025manzano, wu2025openuni] are typically built upon powerful large language models (LLMs) [zhao2023survey] and extend next-token prediction [chen2024next] to image generation: the LLM generates visual tokens, and a VQ-VAE [van2017neural] or Diffusion model serves as a detokenizer. Moreover, Transfusion [zhou2024transfusion] and Show-O [xie2024show] demonstrate that bidirectional image diffusion can be combined with autoregressive text prediction within the same framework. D-DiT [li2024dual] achieves both Text-to-Image (T2I) and Image-to-Text (I2T) tasks using an end-to-end diffusion model. A crucial question concerning unified multimodal models is whether their understanding and generation capabilities can mutually enhance each other. Some studies [tong2024metamorph, wu2024liquid] have provided evidence supporting this phenomenon. However, in contrast to the comprehensive benchmarks for multimodal understanding, T2I benchmarks are often relatively simple, lacking an in-depth examination of complex semantic understanding and world knowledge reasoning, making it difficult to fully prove. Consequently, a new benchmark is imperative to specifically probe these emerging, knowledge-driven generative capabilities.
@@ 2.3 Text to Image Evaluation
Despite the Fréchet Inception Distance (FID) [heusel2017gans] being one of the most widely adopted metrics for evaluating the quality of generated images, it falls short in assessing text-image consistency, thus failing to comprehensively measure the capabilities of text-to-image models. To address this deficiency, researchers have introduced a series of more sophisticated and challenging benchmarks [sim2024evaluating, lin2404evaluating, wu2024conceptmix] and evaluation metrics [hessel2021clipscore, li2024genai, lin2404evaluating, jin2025srum]. For instance, DPG-Bench [hu2024equip] focuses on evaluating models’ ability in dense prompt following. T2I-CompBench [huang2023t2i] provides a benchmark suite for evaluating compositional generation, where prompts typically combine multiple distinct attributes. Furthermore, GenEval [ghosh2024geneval] serves as an object-centric evaluation framework specifically designed to assess compositional attributes of images, such as object co-occurrence, position, number, and color. However, the prompts used in these benchmarks are mostly straightforward, primarily examining models’ ability to follow compositional instructions for generation. Recently, some studies have shifted focus towards evaluating the application of specific types of knowledge in T2I models, such as physical reasoning in PhyBench [meng2024phybench] and broad common-sense knowledge in Commonsense-T2I [fu2024commonsense]. Nevertheless, these emerging benchmarks are still very limited in their research scope and the quantity of evaluations.
@@ 3.2 Discovering the Deep Visual Language Alignment Bottlenecks of Traditional Evaluation
Existing metrics such as CLIP-Score [hessel2021clipscore] and VQA-Score [li2024genai] offer partial solutions for evaluating text-to-image alignment, but both fall short when faced with complex, knowledge-intensive prompts. CLIP-Score, constrained by CLIP’s bag-of-words limitation, struggles with compositional semantics and relational reasoning. VQA-Score improves on this by leveraging likelihood estimation via question answering, but still lacks sensitivity to implicit visual understanding and world knowledge grounding.To address these limitations, we propose a multi-faceted evaluation protocol to rigorously assess the quality of generated images, focusing on four key aspects: Consistency, Realism, Aesthetic Quality, and a composite metric, WiScore.
@@
$\begin{gathered}\frac{\alpha_{1}\times\text{Consistency}+\alpha_{2}\times\text{Realism}+\alpha_{3}\times\text{Aesthetic Quality}}{2},\\ \text{subject to }\alpha_{1}+\alpha_{2}+\alpha_{3}=1.\end{gathered}$ (1)
In this paper, we set $\alpha_{1}=0.7$, $\alpha_{2}=0.2$, and $\alpha_{3}=0.1$. This configuration not only prioritizes Consistency, reflecting the importance of accurately representing the prompt’s intended objects and their relationships, but also incorporates Realism and Aesthetic Quality to ensure overall image quality. A higher WiScore indicates superior performance in accurately depicting objects and concepts based on world knowledge. The detailed scoring criteria is provided in Appendix B. In the evaluation, we employ GPT-4o-2024-05-13, a powerful multimodal large language model, as the evaluator to assess the performance of T2I models. A carefully designed and rigorous scoring protocol is adopted to ensure consistency and reliability, with full implementation details provided in the Appendix C. As shown in Figure 4, among all automated metrics, WiScore exhibits the highest agreement with human judgments. The detailed comparison process can be found in the Appendix D.
Figure 4: Evaluation results of different metrics, where WiScore most closely aligns with human assessment.
@@ 4.1 Experiment Settings
We evaluate 20 T2I models, including 10 dedicated T2I models and 10 unified multimodal models. The dedicated T2I models are: stable-diffusion-v1-5 [rombach2022high], stable-diffusion-2-1 [rombach2022high], stable-diffusion-xl-base-0.9 [podell2023sdxl], stable-diffusion-3-medium [esser2024scaling], stable-diffusion-3.5-medium [esser2024scaling], stable-diffusion-3.5-large [esser2024scaling], playground-v2.5-1024px-aesthetic [li2024playground], PixArt-XL-2-1024-MS [chen2023pixartalpha], FLUX.1-dev [flux2024], and FLUX.1-schnell [flux2024]. For unified multimodal models, we selected 12 models representing three mainstream architectural paradigms for evaluation. These include: Autoregressive (AR) Paradigm models, such as Janus-Pro-7B [chen2025janus], Janus-Pro-1B [chen2025janus], Janus-1.3B [wu2024janus], vila-u-7b-256 [wu2024vila], and Emu3 [wang2024emu3]; Cascaded AR-Diffusion models (using the AR model’s last hidden state to pass condition), including Qwen-Image [wu2025qwen], UniWorld [lin2025uniworld], BLIP3o [chen2025blip3], and MetaQuery [pan2025transfer]; and the Parallel AR-Diffusion model (connecting generation and understanding via shared attention), BAGEL [deng2025emerging]. For image generation, we use the official default configurations of each model and fix the random seed to ensure reproducibility. For short, we use SD to denote stable-diffusion, playground-v2.5 for playground-v2.5-1024px-aesthetic, and PixArt-Alpha for PixArt-XL-2-1024-MS.
@@ 4.2 Main Results
[…]
@@
[…]
Parallel Autoregressive & Diffusion
BAGEL 0.44 0.55 0.68 0.44 0.60 0.39 0.52
BAGEL+CoT 0.76 0.69 0.75 0.65 0.75 0.58 0.70
Cascaded Autoregressive & Diffusion
Qwen-Image 0.62 0.63 0.77 0.57 0.75 0.40 0.62
UniWorld-V1 0.53 0.55 0.73 0.45 0.59 0.41 0.55
@@
Understanding-Generation Gap in T2I Tasks.subsection 4.2 presents the WiScore of various models on the WISE benchmark, designed to evaluate the extent to which models leverage world knowledge in image generation. Overall, the majority of models fail to achieve a satisfactory score (>0.6>0.6), indicating significant deficiencies in the ability of current models to leverage complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching) for image generation. This limitation is particularly evident in dedicated T2I models. However, when examining the overall model architectures, we find that the Parallel Autoregressive & Diffusion and Cascaded Autoregressive & Diffusion paradigms within unified MLLMs outperform dedicated T2I models, achieving the best results. This suggests that these two architectural approaches are most effective at eliciting the model’s internal knowledge, demonstrating the advantages and potential of unified models. Conversely, the Autoregressive (AR) models exhibit generally poor performance, highlighting the difficulty of the AR paradigm in integrating the model’s internal priors into the image generation process.
Notably, when BAGEL utilizes Chain of Thought (CoT) to assist its visual generation, the model achieves the optimal performance on our benchmark. This indicates that CoT is an effective mechanism for eliciting model knowledge, and suggests that training with CoT is an effective strategy for enhancing the application of internal knowledge during the generative process.
@@
[…]
Parallel Autoregressive & Diffusion
BAGEL 0.68 +0.24 0.75 +0.20 0.75 +0.07 0.71 +0.27 0.81 +0.21 0.81 +0.42 0.73 +0.21
Cascaded Autoregressive & Diffusion
Qwen-Image 0.91 +0.29 0.82 +0.19 0.92 +0.15 0.88 +0.31 0.91 +0.16 0.80 +0.40 0.88 +0.26
UniWorld-V1 0.81 +0.28 0.75 +0.20 0.84 +0.11 0.71 +0.26 0.76 +0.17 0.69 +0.28 0.78 +0.23
BLIP3o-8B 0.80 +0.31 0.70 +0.19 0.75 +0.12 0.80 +0.26 0.76 +0.13 0.73 +0.36 0.76 +0.24
Autoregressive
We conducted an additional experiment to further highlight the limitations of current T2I models in processing world knowledge and explore possible solutions. We rewrote the prompts in the WISE benchmark using GPT-4o-2024-05-13, transforming them from complex, knowledge-requiring prompts to straightforward prompts. For example, the original prompt "Common plants for Mother’s Day" was rewritten to "Carnation". The detailed instructions for prompt rewriting are detailed in Appendix E. The same WiScore evaluation method was used for this experiment, and the results are presented in  subsection 4.3.
Nearly all models demonstrated significant performance improvements when evaluated with the rewritten prompts. The top-performing model, Qwen-Image, achieved a remarkable overall WiScore of 0.88. This suggests that a significant portion of the difficulty in the original WISE benchmark stemmed from the models’ inability to effectively parse and contextualize complex, knowledge-intensive prompts. By simplifying the prompts, we effectively removed this comprehension bottleneck, allowing the models to better showcase their inherent generative capabilities.
Notably, the models with the largest overall improvements tended to be those with lower initial scores on the original benchmark. For example, Janus-Pro-1B demonstrated an overall WiScore improvement of +0.34 after rewriting, highlighting the sensitivity of weaker models to prompt complexity. Furthermore, it is noteworthy that BAGEL’s performance on rewritten prompts (0.73 Overall WiScore) is highly comparable to that of BAGEL+CoT on the original prompts (0.70 Overall WiScore). This observation implies that CoT serves a function similar to prompt simplification, acting as an effective mechanism to elicit and integrate the model’s internal knowledge during the generation process.